# ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling

### v1

@@ Abstract
The development of aerial holistic scene understanding algorithms is hindered by the scarcity of comprehensive datasets that enable both semantic and geometric reconstruction. While synthetic datasets offer an alternative, existing options exhibit task-specific limitations, unrealistic scene compositions, and rendering artifacts that compromise real-world applicability. We introduce ClaraVid, a synthetic aerial dataset specifically designed to overcome these limitations. Comprising 16,917 high-resolution images captured at 4032×3024 from multiple viewpoints across diverse landscapes, ClaraVid provides dense depth maps, panoptic segmentation, sparse point clouds, and dynamic object masks, while mitigating common rendering artifacts. To further advance neural reconstruction, we introduce the Delentropic Scene Profile (DSP), a novel complexity metric derived from differential entropy analysis, designed to quantitatively assess scene difficulty and inform reconstruction tasks. Utilizing DSP, we systematically benchmark neural reconstruction methods, uncovering a consistent, measurable correlation between scene complexity and reconstruction accuracy. Empirical results indicate that higher delentropy strongly correlates with increased reconstruction errors, validating DSP as a reliable complexity prior. Currently under review, upon acceptance the data and code will be available at rdbch.github.io/ClaraVid.
@@ 1 Introduction
Semantic and geometric scene understanding derived from low-altitude uncrewed aerial vehicle (UAV) imagery has become increasingly essential for applications such as disaster response  [1], urban mapping [2], and infrastructure inspection [3], where accurate and high-fidelity environmental perception directly informs critical decision-making processes. Recent advancements in neural reconstruction techniques, notably Neural Radiance Fields (NeRFs)[4, 5, 6] and Gaussian Splatting approaches[7, 8, 9], have demonstrated significant potential in enabling scene-level querying and novel-view synthesis, offering a transformative approach to reconstructing complex 3D environments from sparse 2D observations. However, evaluating holistic scene mapping in urban environments—encompassing both semantic and geometric aspects—remains constrained by the limitations of existing datasets. Real-world benchmarks providing comprehensive multi-modal ground truth for the joint tasks are inexistent for urban scenarios, and synthetic alternatives suffer from critical shortcomings: they often lack key sensing modalities such as dense semantic segmentation [10, 11], are not collected for scene mapping applications [12, 13], or are captured from restricted viewpoints that hinder a full assessment of reconstruction quality and generalization [14].
@@
In this work, we introduce ClaraVid, a synthetic aerial dataset crafted to support the training and evaluation of holistic scene representations while also enabling lower-level perception tasks. ClaraVid employs a mapping-centric approach, emulating UAV mapping missions to reflect real-world aerial survey workflows. Spanning 5 distinct environments across 8 missions, the dataset provides high-resolution imagery at 4032 × 3024 pixels, collected simultaneously from 3 different viewpoints, accompanied by multimodal annotations including dense depth maps, panoptic segmentation for buildings and actors, and dynamic object masks to precisely delineate moving entities—a distinctive enhancement over prior synthetic datasets. Furthermore, it includes multi-resolution multi-modal scene-level point clouds approximating LiDAR scans, broadening its utility for geometric analysis. Data quality is ensured through an optimized rendering pipeline that addresses challenges such as foliage culling and level-of-detail inconsistencies, resulting in enhanced fidelity for scene reconstruction and semantic segmentation. The dataset is complemented by a benchmarking framework designed to assess semantic and geometric reconstruction accuracy, while also supporting frame-level tasks such as semantic segmentation. Relevant image samples are presented in ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling.
@@
A major challenge in UAV-based scene reconstruction is quantifying scene complexity in a way that correlates with the difficulty of downstream reconstruction tasks. Currently the base solution is based on an heuristic assessments rather than formalized metrics, making it difficult to systematically predict the impact of scene structure on the reconstruction. Delentropy[15] , an entropy measure derived from gradient field analysis, has previously been employed to analyze information content in pixel-level tasks [16, 17, 18]. In this work, we adapt delentropy for scene profiling, and propose a novel Delentropic Scene Profile, applied in the context of neural reconstruction, leveraging its ability to capture the spatial organization of scene features through second-order entropy analysis. We empirically demonstrate a strong correlation between the proposed metric and the final reconstruction error across multiple approaches.
@@
To summarize, our paper introduces the following primary contributions to advance UAV-based holistic scene understanding: (1) We present ClaraVid, a novel synthetic aerial dataset specifically designed for holistic scene reconstruction, comprising high-resolution imagery, multi-modal annotations (dense depth maps, panoptic segmentation, sparse point clouds, and dynamic object masks), and realistic rendering tailored to aerial mapping applications. (2) We introduce the Delentropic Scene Profile (DSP), a delentropy-based metric for quantitatively assessing scene complexity. (3) We empirically validate DSP’s utility, demonstrating strong correlations between scene complexity and reconstruction performance, offering a principled framework for guiding method selection and optimizing neural reconstruction pipelines.
@@
Table 1: Comparison of related datasets from aerial and automotive domains relevant for holistic scene representation.
@@
Dataset Year Type Camera Calib. Seg. Depth Scene PCL Dynamic Mask For Mapping Viewpoints Render Issues Image Size Samples
@@
UrbanScene3D [11] 2022 S/R Instance* RGB/Ins 1 - 8192x5460 128k
Ommo [25] 2023 R RGB 1 - 4096x2160 14.7k
Matrix City [10] 2023 S RGB 1 Low* 1920x1080 410k
@@
ClaraVid (ours) - S Panoptic RGB/Sem/Ins 3 Low* 4032x3024 16.3k
@@ 2 Related work
Aerial datasets The progress in semantic and geometric scene understanding has been largely driven by multi-modal automotive datasets, real [26, 19, 20] and synthetic [27, 28, 21]. These datasets provide strong benchmarks that integrate LiDAR, stereo imagery, and detailed annotations for tasks enabling autonomous navigation. In contrast, real-world aerial datasets, are often small in scale and fragmented in scope: semantic segmetation[24, 29, 22, 30], depth estimation [31, 32], object tracking [33], etc. Synthetic aerial datasets, created using graphics engines, offer a scalable alternative with rich semantic content. However, they come with significant limitations. Their resolution is typically limited to 2K [13], whereas real datasets can reach 4K [24, 22, 29]. Rendering artifacts, such as issues with foliage rendering [34] and inconsistent levels of detail [13, 12], reduce their realism. Additionally, key modalities like semantic segmentation are frequently missing [21, 10], making them less effective for mapping applications. Claravid addresses these limitations directly by providing a dataset specifically designed for mapping tasks. It includes multi-view imagery, panoptic annotations, and dense depth maps, offering a high-fidelity benchmark targeted towards real-world applications. A summarized overview of the presented datasets is provided in Table 1.
@@
Neural scene representation of urban scenarios from aerial perspective Recent advancements in neural scene representations have established robust frameworks for modeling complex, real-world environments. NeRFs, introduced by Mildenhall etal[4], have proven highly effective across a range of tasks–where further extensions–shown promising results from novel view synthesis [35, 36] to surface reconstruction [37, 38, 39, 40] and panoptic reconstruction [41, 42]. In parallel, Gaussian Splatting [7] has emerged as a particle-based alternative that bypasses the computational overhead of neural networks, thereby enabling significantly accelerated training and rendering for real-time, large-scale scene representations [7, 43, 8, 44]. However, many of these approaches are relying on prior data for semantic information that can be obtained typically from GT data or as pseudolabels. Finally, while both paradigms have been applied to urban scene reconstruction from aerial perspectives [45, 46], they largely omit semantic information—primarily due to the absence of comprehensive datasets. We introduce our dataset specifically designed to benchmark these diverse methodologies within a unified framework that integrates all relevant modalities from a low aerial viewpoint.
@@
Dataset complexity through the information-theory lens Differential entropy has emerged as a robust information-theoretic tool for quantifying image complexity, building on Shannon’s definition of entropy[47] as a measure of uncertainty in data. Larkin’s formulation of delentropy [15] exemplifies this approach, extending traditional entropy to incorporate spatial structure by using a gradient-based joint density (deldensity) that captures pixel co-occurrence and image detail. This enables a more comprehensive complexity measure than conventional metrics: whereas basic Shannon entropy ignores spatial relationships among pixels and GLCM-based texture entropy[48] captures only limited local patterns, delntropy integrates both local and global features of an image’s intensity distribution. Rahane et al.[16] compared large-scale image datasets using entropy distributions (pixel-level, texture, and delentropy), finding that higher-order measures like delentropy align well with dataset learning difficulty and correlate with known variations in model performance for the semantic segmentation task. Similarly, Cagas et al. [18] observed that a dataset’s complexity distribution—quantified via delentropy—correlates with generative model fidelity, informing how increasing image complexity can demand larger training sets for GANs[49]. We leverage delentropy to construct a scene-level complexity measure that serves as a proxy for assessing reconstruction difficulty in aerial missions.
@@ 3.1 Theoretical background
Image complexity can be defined as the variety of features and details within an image. Traditional entropy is a foundational abstraction in information theory introduced by Shannon[47]. It is used as a measure of uncertainty in data, specifically the variation in the distribution of pixel intensities of an image in grayscale format. To formalize delentropy mathematically, we start with its foundation in the gradient field of an image. Given an image $I$, the gradient components are computed as the partial derivatives:
@@
$p_{i,j}=\frac{1}{4XY}\sum_{n=-N}^{N-1}\sum_{m=-M}^{M-1}\delta_{i,f_{x}(m,n)}\delta_{j,f_{y}(m,n)}$ (2)
where $p_{ij}$ represents the joint probability of observing the gradient values (i,j) at any pixel, $N,M$ define the spatial window size over which the gradient distribution is computed and $X,Y$, the image width and height. Further the sum over all non-zero gradient value pairs is normalized such that it represents a probability density function $\sum_{i,j}p_{ij}=1$. Delentropy[15], denoted as $H_{\text{del}}$, is then defined as the joint entropy of this distribution. If $p_{ij}$ is the probability of $f_{x}$ and $f_{y}$ falling into the $i$-th and $j$-th bins, respectively, the discrete form is:
@@
such that $I$ and $J$ represent the number of bins (discrete cells) in the 2D distribution, while the factor $\tfrac{1}{2}$ is derived from  [51] and is maintained in our case to preserve the scale consistency without significantly affecting the metrics. Further we will denote $H_{del}(I)$ the delentropy of image $I$.
@@ 3.2 Delentropic Scene Profile
The delentropy measure, $H_{\text{del}}$, quantifies structural complexity for an individual image. To extend this to a collection of images at scene level, we introduce the novel concept of Delentropic Scene Profile ($DSP_{S}$), a quantitative framework capturing the complexity of a multi-view scene and its implications for reconstruction accuracy. A scene $S$ is a finite set of images, $S=\{I_{1},I_{2},\ldots,I_{N}\}$, where each $I_{k},k=1,\ldots,N$ corresponds to a unique viewpoint of a shared environment, collectively encoding its 3D structure. The delentropy set of $S$ is defined as $\{H_{\text{del},k}\}_{k=1}^{N}$, with $H_{\text{del},k}=H_{\text{del}}(I_{k})$ computed over the grayscale pixels relevant to reconstruction, assigning zero informational content to excluded regions (e.g., masked dynamic objects) to preserve measure integrity.
The Delentropic Scene Profile, $\text{DSP}_{S}$, is the statistical distribution of $\{H_{\text{del},k}\}$, explicitly aproximated to a truncated Beta distribution parameterized by $(\alpha,\beta,a,b)$:
@@
In the absence of closed-form solutions, we first estimate $\alpha$ and $\beta$ using the Method of Moments (MoM), with the sample moments appropriately adjusted for the support $[a,b]$. These MoM estimates then serve as initial values for the Newton-Raphson algorithm, which iteratively refines the parameters to maximize the likelihood function.
@@
3.3 Motivation for Beta Distribution
The choice of distributional model for $\text{DSP}_{S}$ was guided by the analysis of delentropy profiles of diverse datasets—ClaraVid, Skydrone[12], Skyscapes[30], UAVid[24], 3D Matrix City [10], Mill19 [46], and UrbanScene3D [11]—observing varied distributional forms contingent on the reconstruction context. A Gaussian distribution, while apt for uniformly complex scenes, fails to capture skewed or long-tailed distributions prevalent in heterogeneous environments (for example where we have a body of water or sky regions). Alternatives such as the Pareto or Gamma distributions accommodate skewness but lack flexibility for U-shaped profiles, which emerge in scenarios with pronounced bimodal complexity (e.g., a field with ocassional high-detail trees). The Beta distribution, however, adeptly models this spectrum—uniform, skewed, and U-shaped—owing to its shape parameters $\alpha$ and $\beta$, requiring only four parameters (including bounds $a$ and $b$) for a parsimonious yet expressive fit. This simplicity affords intuitive interpretation: $\alpha$ and $\beta$ directly govern the density’s form, while $a$ and $b$ anchor it to the empirical delentropy range. Multimodal distributions were eschewed, as their complexity—necessitating mixture models with additional parameters—hampers interpretability, risks overfitting, particularly given the finite sample sizes of $\{H_{\text{del},k}\}$ and did not show any improvements as compared with the best one. The Beta distribution emerges as the best choice, balancing flexibility, robustness, and theoretical clarity.

### v2

@@ Abstract
The development of aerial holistic scene understanding algorithms is hindered by the scarcity of comprehensive datasets that enable both semantic and geometric reconstruction. While synthetic datasets offer an alternative, existing options exhibit task-specific limitations, unrealistic scene compositions, and rendering artifacts that compromise real-world applicability. We introduce ClaraVid, a synthetic aerial dataset specifically designed to overcome these limitations. Comprising 16,917 high-resolution images captured at 4032×3024 from multiple viewpoints across diverse landscapes, ClaraVid provides dense depth maps, panoptic segmentation, sparse point clouds, and dynamic object masks, while mitigating common rendering artifacts. To further advance neural reconstruction, we introduce the Delentropic Scene Profile (DSP), a novel complexity metric derived from differential entropy analysis, designed to quantitatively assess scene difficulty and inform reconstruction tasks. Utilizing DSP, we systematically benchmark neural reconstruction methods, uncovering a consistent, measurable correlation between scene complexity and reconstruction accuracy. Empirical results indicate that higher delentropy strongly correlates with increased reconstruction errors, validating DSP as a reliable complexity prior. The data and code are available on the project page: rdbch.github.com/claravid.
@@ 1 Introduction
Semantic and geometric scene understanding derived from low-altitude uncrewed aerial vehicle (UAV) imagery has become increasingly essential for applications such as disaster response  [1], urban mapping [2], and infrastructure inspection [3], where accurate and high-fidelity environmental perception directly informs critical decision-making processes. Recent advancements in neural reconstruction techniques, notably Neural Radiance Fields (NeRFs)[4, 5, 6] and Gaussian Splatting (GSplat) approaches[7, 8, 9], have demonstrated significant potential in enabling scene-level querying and novel-view synthesis, offering a transformative approach to reconstructing complex 3D environments from sparse 2D observations. However, evaluating such holistic reconstruction of urban environments—encompassing both semantic and geometric aspects—remains constrained by the limitations of existing datasets. Real-world benchmarks providing comprehensive multi-modal ground truth for the joint tasks are inexistent for urban scenarios, and synthetic alternatives suffer from critical shortcomings: they often lack key sensing modalities such as dense semantic or instance segmentation [10, 11], are not collected for scene mapping applications [12, 13], or are captured from restricted viewpoints that hinder a full assessment of reconstruction quality and generalization [14].
@@
In this work, we introduce ClaraVid, a synthetic aerial dataset crafted to support the training and evaluation of holistic scene reconstruction while also enabling lower-level perception tasks. ClaraVid employs a mapping-centric approach, emulating UAV mapping missions to reflect real-world aerial survey workflows. Spanning 5 distinct environments across 8 missions, the dataset provides high-resolution imagery, collected simultaneously from 3 different viewpoints, accompanied by multimodal annotations: dense depth maps, panoptic segmentation and dynamic object masks to delineate moving entities–a distinctive enhancement over prior synthetic datasets. Furthermore, it includes multi-resolution multi-modal scene-level point clouds approximating LiDAR scans, broadening its utility for geometric analysis. Data quality is ensured through an optimized rendering pipeline that addresses challenges such as foliage culling and level-of-detail inconsistencies, resulting in enhanced fidelity for scene reconstruction and semantic segmentation. Notably, Claravid systematically spans structural complexities–from homogeneous areas to highly detailed urban environments–quantified via scene complexity profile. The dataset is complemented by a benchmarking framework designed to assess semantic and geometric reconstruction accuracy at various viewpoints, while also supporting frame-level tasks (depth estimation, panoptic segmentation, etc). An overview is presented in ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling.
@@
A major challenge in UAV-based neural scene reconstruction is quantifying scene complexity in a way that correlates with the difficulty of downstream reconstruction tasks (e.g., novel view synthesis) only from 2D measurements. Currently, the base solution relies on heuristic assessments rather than formalized metrics, making it difficult to systematically predict the impact of scene structure on the reconstruction. Delentropy[15], an entropy measure derived from gradient field analysis, has previously been employed to analyze information content in pixel-level tasks [16, 17, 18]. In this work, we adapt delentropy to quantify scene complexity in neural reconstruction and introduce the Delentropic Scene Profile as a descriptor of scene-level complexity. We empirically demonstrate a strong correlation between the proposed metric and the final error across multiple reconstruction approaches.
@@
To summarize, our paper introduces the following contributions to advance UAV-based holistic scene understanding: (1) We present ClaraVid, a novel synthetic aerial dataset specifically designed for holistic neural scene reconstruction, comprising high-resolution imagery, multi-modal annotations, and realistic rendering tailored for neural scene reconstruction tasks from an aerial perspective. (2) We introduce the Delentropic Scene Profile, a delentropy-based metric for quantitatively assessing scene complexity only from 2D measurements. (3) We empirically demonstrate the utility of both the dataset and DSP through extensive experiments.
@@
Table 1: Related Aerial Datasets–Comparison of datasets from aerial and automotive domains relevant for holistic scene representation.
@@
Dataset Year Type* Camera Calib. Seg. Depth Scene PCL Dynamic Mask For Mapping Viewpoints Render Issues Image Size Samples
@@
UrbanScene3D [11] 2022 S/R Instance RGB/Ins 1 - 8192x5460 128k
Dronescapes [26] 2023 R Semantic 1 - 3840x2160 16.9k
Matrix City [10] 2023 S RGB 1 Low 1920x1080 410k
@@
ClaraVid (ours) S Panoptic RGB/Sem/Ins 3 Low 4032x3024 16.9k
*S/R - Synthetic / Real
@@ 2 Related work
Aerial datasets The progress in semantic and geometric scene understanding of urban scenes has been largely driven by multi-modal automotive datasets, real [27, 19, 20] and synthetic [28, 29, 21]. These datasets provide strong benchmarks that integrate LiDAR, multi-view imagery, and detailed annotations for tasks enabling autonomous navigation. In contrast, real-world aerial datasets, are often small in scale and fragmented in scope: semantic segmentation [22, 24, 25, 30, 31], depth estimation [32, 33], object tracking [34, 35], LiDAR segmentation [36], etc. While some works [26] exist that integrate both semantic and geometric information, none of them are targeted towards mapping applications. Synthetic aerial datasets, created using graphics engines, offer a scalable alternative with rich semantic content. However, they come with significant limitations. Their resolution is typically limited to 2K [13], whereas real datasets can reach 4K [25, 22, 31]. Rendering artifacts, such as issues with foliage rendering [37] and inconsistent levels of detail [13, 12], reduce their realism. Additionally, key modalities like semantic segmentation are frequently missing [21, 10, 38], making them less effective for semantic mapping applications. ClaraVid addresses these limitations directly by providing a dataset specifically designed for mapping tasks. It includes multi-view imagery, panoptic annotations, and dense depth maps, offering a high-fidelity benchmark targeted towards real-world applications. A summary of the presented datasets is provided in Table 1.
@@
Neural scene representation of urban scenarios from aerial perspective Recent advancements in neural scene representations have established robust frameworks for modeling complex, real-world environments. NeRFs, introduced by Mildenhall et al.[4], have proven highly effective across a range of tasks–where further extensions shown promising results from novel view synthesis [39, 40] to surface reconstruction [41, 42, 43, 44] and panoptic reconstruction [45, 46]. In parallel, Gaussian Splatting [7] has emerged as a particle-based alternative that bypasses the computational overhead of neural networks, thereby enabling significantly accelerated training and rendering for real-time, large-scale scene representations [7, 47, 8, 48]. However, many of these approaches are relying on prior data for semantic information that can typically be obtained from GT data or as pseudolabels. Finally, while both paradigms have been applied to urban scene reconstruction from aerial perspectives [49, 50], they largely omit semantic information, primarily due to the absence of comprehensive datasets. We introduce our dataset specifically designed to benchmark these diverse methodologies within a unified framework that integrates all relevant modalities from a low aerial viewpoint.
@@
Dataset complexity through the information-theory lens Differential entropy has emerged as a robust information-theoretic tool for quantifying image complexity, building on Shannon’s definition of entropy[51] as a measure of uncertainty in data. Larkin’s formulation of delentropy [15] exemplifies this approach, extending traditional entropy to incorporate spatial structure by using a gradient-based joint density (deldensity) that captures pixel co-occurrence and image detail. This enables a more comprehensive complexity measure than conventional metrics: whereas basic Shannon entropy ignores spatial relationships among pixels and GLCM-based texture entropy[52] captures only limited local patterns, delentropy integrates both local and global features of an image’s intensity distribution. Rahane et al.[16] compared large-scale image datasets using entropy distributions (pixel-level, texture, and delentropy), finding that higher-order measures like delentropy align well with dataset learning difficulty and correlate with known variations in model performance for the semantic segmentation task. Similarly, Cagas et al. [18] observed that a dataset’s complexity distribution, quantified via delentropy, correlates with generative model fidelity, informing how increasing image complexity can demand larger training sets for GANs[53]. We leverage delentropy to construct a scene-level complexity measure that serves as a proxy for assessing reconstruction difficulty in aerial missions only from 2D images.
@@ 3 Scene complexity through delentropy
In UAV-based scene reconstruction, we focus on predicting reconstruction complexity directly from 2D aerial imagery, targeting novel view synthesis without access to any reconstructed geometry. This formulation departs from prior work[54] that computes complexity during or after 3D reconstruction. Instead, our approach estimates complexity a priori from images alone. We restrict the problem to structured mapping scenarios where data acquisition follows grid-based trajectories at discrete altitudes, with bounded spatial extents and minimized peripheral content (e.g., distant backgrounds). These constraints reflect real-world UAV mission protocols and provide a stable foundation for inferring scene difficulty from image characteristics alone. Within this context, we formalize the task of image-based complexity estimation and lay the groundwork for linking it to downstream reconstruction error through a principled framework.
@@ 3.1 Theoretical background
We interpret image complexity from an information-theoretic perspective, as the variability of spatial structure encoded within a single view. Classical formulations, such as Shannon’s pixel entropy, capture only the intensity distributions, disregarding spatial dependencies. To generalize spatial entropy in a principled manner, we adopt Larkin’s delentropy[15], a reinterpretation of Shannon’s second-order entropy, where entropy is computed over the gradient field. Delentropy can be formalized as follows, given an image $I$, the gradient components are computed as the partial derivatives:
@@
$p_{ij}=\frac{1}{4XY}\sum_{n=-N}^{N-1}\sum_{m=-M}^{M-1}\delta_{i,f_{x}(m,n)}\delta_{j,f_{y}(m,n)}$ (2)
where $p_{ij}$ represents the joint probability of observing the gradient values (i,j) at any pixel, $N,M$ define the spatial window size over which the gradient distribution is computed and $X,Y$, the image width and height. Delentropy, denoted as $H_{\text{del}}$, is then defined as the joint entropy of this distribution. If $p_{ij}$ is the probability of $f_{x}$ and $f_{y}$ falling into the $i$-th and $j$-th bins, respectively, the discrete form is:
@@
such that $I$ and $J$ represent the number of bins in the 2D distribution, while the factor $\tfrac{1}{2}$ is derived from  [56] and maintained for scale consistency.
This formulation assumes that the spatial distribution of image gradients reflects the underlying structural complexity of the scene. Regions with rich texture, repeated patterns, or fine-grained variations yield broader gradient distributions and thus higher delentropy. Conversely, flat or uniform areas lead to concentrated gradient densities and lower delentropy. This provides a compact measure of spatial complexity directly observable from a single image. A graphical example for different complexity levels is provided in Figure 2.
@@ 3.2 Delentropic Scene Profile
We define a scene complexity profile as a statistical characterization of the representational difficulty associated with a finite set of images $\{I_{k}\}_{k=1}^{N}$ depicting a shared 3D environment. Each image $I_{k}$ is assigned a scalar complexity measure $C_{k}\in\mathbb{R}_{\geq 0}$, and the resulting collection $\{C_{k}\}_{k=1}^{N}$ induces a distribution that captures the variability of structural content across views. Under structured acquisition protocols (e.g., UAV trajectories with regular grid-based sampling and bounded spatial extent), each image encodes a spatially distinct yet complementary subset of the environment. In such settings, the aggregated complexity distribution offers a principled descriptor of scene-level heterogeneity and serves as a proxy for the intrinsic challenge posed to reconstruction algorithms.
To instantiate this framework, we adopt delentropy $H_{\mathrm{del}}$ as the underlying image complexity measure. For a scene $S=\{I_{1},\dots,I_{N}\}$, we define the corresponding Delentropic Scene Profile (DSP) as the empirical distribution of per-image delentropy values $\{H_{\mathrm{del},k}\}_{k=1}^{N}$, with $H_{\mathrm{del},k}=H_{\mathrm{del}}(I_{k})$. These values are computed over regions deemed relevant for reconstruction (e.g., excluding masked dynamic content). We model this distribution via a truncated Beta distribution parameterized by $(\alpha,\beta,a,b)$: